# Lookahead: An Inference Acceleration Framework for Large Language Model with Lossless Generation Accuracy

### v2

@@ Abstract.
As Large Language Models (LLMs) have made significant advancements across various tasks, such as question answering, translation, text summarization, and dialogue systems, the need for accuracy in information becomes crucial, especially for serious financial products serving billions of users like Alipay. To address this, Alipay has developed a Retrieval-Augmented Generation (RAG) system that grounds LLMs on the most accurate and up-to-date information. However, for a real-world product serving millions of users, the inference speed of LLMs becomes a critical factor compared to a mere experimental model.
Hence, this paper presents a generic framework for accelerating the inference process, resulting in a substantial increase in speed and cost reduction for our RAG system, with lossless generation accuracy. In the traditional inference process, each token is generated sequentially by the LLM, leading to a time consumption proportional to the number of generated tokens. To enhance this process, our framework, named lookahead, introduces a multi-branch strategy. Instead of generating a single token at a time, we propose a Trie-based Retrieval (TR) process that enables the generation of multiple branches simultaneously, each of which is a sequence of tokens. Subsequently, for each branch, a Verification and Accept (VA) process is performed to identify the longest correct sub-sequence as the final output. Our strategy offers two distinct advantages: (1) it guarantees absolute correctness of the output, avoiding any approximation algorithms, and (2) the worst-case performance of our approach is equivalent to the conventional process. We conduct extensive experiments to demonstrate the significant improvements achieved by applying our inference acceleration framework. Code is avaliable: https://github.com/alipay/PainlessInferenceAcceleration. footnotetext: Preprint. Under review.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: World Wide Web Conference; May 13-17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large language models (LLMs) based on transformer architecture have exhibited remarkable performance across various benchmarks, rendering them highly valuable in numerous industries. However, despite their significant achievements in language-based tasks, LLMs still face challenges in terms of inference latency when employed in generative tasks. This drawback becomes particularly apparent in scenarios where step-wise decoding is implemented. For instance, with the GLM [1] 10B model using a single Nvidia A100 GPU, the inference speed’s theoretical limit through step-wise decoding is merely 101 tokens per second.
To further gain deeper insights into the factors affecting the LLMs’ inference latency, we conduct a comprehensive theoretical analysis focused on a specific instance. It shows that the input-output (IO) and computation consumption time for inferring each token through step-wise decoding is about 10ms and 0.07ms, respectively [^1] Memory bandwidth of A100 is 2039 GB/s, a 10B model with float16 need more than 20 GB IO, the IO time is about 10ms (20/2039), while A100 has 312 TFLOPS for float16 datatype, and forward a token need about 20B FLOPs, the calculation time is about 0.06ms (20/312000). Therefore, rather than the computation consumption time that is related to the hardware’s FLOPs capacity, the IO consumption time that is highly associated with the model’s size and the memory bandwidth is the key that significantly affects the LLMs’ overall inference latency.
@@
Various techniques, such as quantization [2, 3], sparsity [4, 5], pruning [6, 7], distilling [8, 9], and tensor decomposition [10, 11], have been proposed to reduce the LLMs’ size and the IO consumption time for predicting each token in LLMs. However, these techniques have been found to result in a degradation of accuracy. To address the challenge of predicting more tokens within a fixed IO consumption time, non-autoregressive neural machine translation (NAT) [12] and iterative parallel decoding [13] have been introduced and successfully applied in translation tasks. Unfortunately, this approach has shown limited effectiveness in question-answering scenarios.
Recently, decoding with a draft model has become a popular strategy, as discussed in [14, 15]. However, this strategy necessitates additional training efforts or the search for an available draft model. In light of this, a strategy named LLMA that eliminates the need for a draft model has been developed, as presented in [16]. This strategy involves matching text spans from retrieved documents without requiring significant additional efforts. Despite these merits, it has been observed that this strategy performs poorly beyond document-retrieving scenarios.
@@
Table 1 presents a summary of the aforementioned acceleration methods. It is important to acknowledge that these methods only make use of tokens in a single draft manner, thereby not fully leveraging the computational capabilities of GPUs in parallel. In order to maximize the benefits offered by GPUs, we have developed Lookahead, a groundbreaking framework that incorporates a multi-branch strategy. Lookahead utilizes a trie tree to meticulously record the input and output token lists, allowing for the prediction of multiple paths based on the provided context tokens through a parallel Verification and Accept (VA) process. Furthermore, we have implemented an adaptive strategy to optimize the path-retrieving process, effectively striking a balance between memory and computation requirements. It shows that Lookahead proves its superior performance in accelerating LLMs’ inference, compared with the existing state-of-the-art (SOTA) acceleration method.
@@
Methods Training- Accuracy- Full-FLOPs-
free lossless utility
Quantization [2, 3] - $\times$ $\times$
Sparsity [4, 5] - $\times$ $\times$
Distilling [8, 9] $\times$ $\times$ $\times$
Tensor-decomposition [10, 11] - $\times$ $\times$
Early-exit [17, 18] $\times$ $\times$ $\times$
Block decoding [14] $\times$ $\times$
Speculative decoding [15, 19] - $\times$
LLMA [16] $\times$
Lookahead
@@
Moreover, to gain a wide range of applications, we have implemented our framework based on the transformers library of Hugging face [^2] https://huggingface.co/, by extending a generation mode named lookahead generation, which supports the greedy search and sample strategy with batch inference supported. We have also currently applied Lookahead to the most recent LLMs, such as GLM[1], Llama[20], OPT[21], GPT2[22], BLOOM[23], ChatGLM[24], Baichuan[25] and Qwen[26], etc. The aforementioned models can be easily adapted to integrate Lookahead, our well-designed framework, with only minor code modifications of approximately 20 lines. We will soon release our work with open source.
@@
  • We empirically quantify that the main bottleneck of LLM inference is the IO bandwidth, rather than the FLOPs.
@@
  • We elaborately adapt Lookahead to the most recent LLMs and will soon release our work with open source.
@@ 2. Related Work
Recently, several strategies have been proposed and developed to enhance the inference speed of LLMs while maintaining the output quality within an acceptable range. One such strategy is the non-auto-regressive approach, specifically non-auto-regressive translation (NAT) [12], primarily employed in translation tasks [27, 28, 29]. However, it is essential to note that there are significant distinctions between translation tasks and general language model (LLM) scenarios, which may lead to subpar performance when applying the NAT strategy to LLM decoding.
@@
To address this limitation, Huang et al. [28] introduce a layer-wise iterative method wherein each layer utilizes the decoding results and embeddings from the preceding layers. This approach involves training each layer through maximum likelihood estimation to predict the outcomes of subsequent decoding layers. On the other hand, Santilli et al. [13] formalized the standard greedy auto-regressive decoding strategy by employing a parallel Jacobi and Gauss-Seidel fixed-point iteration. It initializes the next tokens using special tokens and performs iterative decoding until convergence. However, all these methods may suffer from the risk of accuracy degeneration, since the manipulated model deviates from its original version.
@@
Recently, there has been a proposal for accuracy-lossless acceleration to enhance the auto-regressive decoding strategy as illustrated in Table 1. In this approach, a block-wise parallel decoding strategy was introduced by [14]. In this strategy, each subsequent token is independently and parallelly predicted as a proposal using an additional transformer model, which consists of a multi-output feed-forward layer and can be fine-tuned or distilled for optimal performance. Then the proposals are directly compared against the output tokens generated by the original decoder. The longest verified tokens are then selected for acceptance as the current step’s output. Despite this strategy’s remarkable acceleration, more efforts are needed to train the extra feed-forward layer, which is essential to ensure that the predictions generated by the transformer model are accurate and reliable. On the other hand, due to the parallel mechanism employed in this strategy, the prediction process solely relies on the input tokens and does not consider the information of the next token, undermining the prediction’s accuracy. Furthermore, frequent failures during the verification process shorten the output length for each decoding step.
@@
To address the aforementioned issue, speculative decoding has been proposed [15, 19], through utilizing a smaller model as a draft predictor. For instance, the Bloom model, which has a size of 7.1B, can be employed as a small model for the larger 176B model. However, obtaining a qualified small model often becomes impractical due to the difficulty of generating similar sentence pieces as the large model. To overcome this challenge, model-free prediction strategies have been introduced to achieve accurate predictions without relying on a specific model. One such strategy is presented in Ge et al. [30], which utilizes an input-guided method that copies content from the input sentence through prefix matching. Another strategy LLMA, proposed by Yang et al. [16], employs a prefix matching approach to retrieve content from either the input sentence or a document database. These model-free strategies are utilized in a greedy decoding manner, ensuring accuracy while achieving performance gains through approximate inference. This inference technique relaxes the verification criterion by accepting predictions within the top-k scores [14]. However, it is worth noting that all the methods mentioned above utilize tokens in a single draft manner, failing to fully utilize the GPUs.
@@ 3.3. Singe-Branch Strategy
Several most recent methodologies [14] [16] have been proposed to generate a sequence of tokens at each decoding step, with the purpose of promising a higher l to accelerate the LLMs’ inference speed V. In these works, according to the prompt and the output tokens at the previous decoding step, a branch of tokens, named single-branch draft, have been obtained through the small model or the document reference, and efficiently validated by running the LLM in a single forward process.
Despite the single-branch strategy’s success in accelerating LLMs’ inference, the generated tokens’ length per decoding step, namely l, cannot be guaranteed as long as we wish. The previous works [14] [16] empirically conclude that there is an upper limit for l and the inference speed V.
@@ 3.4. Discovering GPU’s FLOPs Redundancy
To explore how far EDL can be extended at each decoding step, we conduct a case study through discussing the decoding length’s impact on the overall consuming time of LLMs’ single forward process, whose result can be found in Figure 1. It should be noted that we apply AntGLM-10B (which is trained from scratch with the GLM structure [1, 24]) model with single Nvidia A100 GPU as an example.
Figure 1. Decoding length’s impact on the overall consuming time of LLMs’ single forward process.
In Figure 1, it is evident that as the decoding length increases in the LLM’s single forward process, the time consumption remains relatively constant in the initial stage, given a fixed context length. This observation can be attributed to several factors. Firstly, certain time requirements, such as kernel launching, operator loading, reading, and verification, are fixed and dependent on the GPU bandwidth. These factors are unrelated to the computational complexity or decoding length. Secondly, when the decoding length is relatively short, the resulting small matrix blocks may not take full advantage of CUDA and Tensor cores, indicating that GPU shows its FLOPs redundancy in this scenario. Consequently, the differences in overall time consumption between decoding lengths of 16 and 128 in the LLM’s single forward process is negligible, provided the context length is 64 or 256. This finding supports the concept that the decoding time per decoding step, represented as t(l), remains nearly constant within a specific range of generated token lengths l. However, when the decoding length exceeds 128 while the context length is 64 or 256 in the LLM’s single forward process, the larger matrix block size introduces a more complex calculation logic for the tensor cores. Consequently, the overall inference time shifts to the second stage, leading to a gradual increase in overall inference time.
@@
It can be clearly seen that there is still a ”sufficient gap” between EDLs empirically concluded previously [14] [16] and the CDL illustrated in Figure 1. In this scenario, we come up with two questions:
@@ 4.1. Overview of Lookahead
In contrast to the single-branch strategy such as [16], which only considers one available draft without considering the existence of other drafts, the multi-branch strategy retrieves multiple drafts simultaneously. These drafts are then efficiently decoded and validated in parallel through the Verification and Accept (VA) process, progressively. The VA process then identifies the correct sub-sequence for each draft and retains the longest sub-sequence as the output tokens. Figure 2 presents the progress overview including various draft retrieving and verifying strategies.

### v3

@@ Abstract.
As Large Language Models (LLMs) have made significant advancements across various tasks, such as question answering, translation, text summarization, and dialogue systems, the need for accuracy in information becomes crucial, especially for serious financial products serving billions of users like Alipay. However, for a real-world product serving millions of users, the inference speed of LLMs becomes a critical factor compared to a mere experimental model.
Hence, this paper presents a generic framework for accelerating the inference process, resulting in a substantial increase in speed and cost reduction for our LLM-based scenarios, with lossless generation accuracy. In the traditional inference process, each token is generated sequentially by the LLM, leading to a time consumption proportional to the number of generated tokens. To enhance this process, our framework, named lookahead, introduces a multi-branch strategy. Instead of generating a single token at a time, we propose a Trie-based retrieval and verification mechanism to be able to accept several tokens at a forward step. Our strategy offers two distinct advantages: (1) it guarantees absolute correctness of the output, avoiding any approximation algorithms, and (2) the worst-case performance of our approach is equivalent to the conventional process. We conduct extensive experiments to demonstrate the significant improvements achieved by applying our inference acceleration framework. Our framework is widely deployed in Alipay since April 2023, and obtain remarkable 2.66x to 6.26x speedup. Our code is available at https://github.com/alipay/PainlessInferenceAcceleration.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25-29, 2024; Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Large language models (LLMs) based on transformer architecture have exhibited remarkable performance across various benchmarks, rendering them highly valuable in numerous industries. However, despite their significant achievements in language-based tasks, LLMs still face challenges in terms of inference latency when employed in generative tasks. This drawback becomes particularly apparent in scenarios where step-wise decoding is implemented.
To further gain deeper insights into the factors affecting the LLMs’ inference latency, we conduct a comprehensive theoretical analysis focused on a specific instance. Generating a single token with a 10B LLM (without sparse activation) necessitates reading approximately 20GB of memory and performing approximately 20G FLOPs. Although the computational load and memory access associated with the attention mechanism scale quadratically and linearly with sequence length, respectively, their impact can be considered negligible when the sequence length is less than half of the hidden dimension. For instance, in a GLM-10B model with a prompt length of 512, the memory required to read weights is 20.27GB, with additional memory for inputs and outputs totaling about 1.63GB. The computational FLOPs for matrix multiplication involving these weights are around 20.27T, whereas the combined FLOPs for attention and activation functions are approximately 0.405T. Given that an Nvidia A100 GPU provides a bandwidth of 2039 GB/s and a computational capacity of 312T FLOPs, the input/output (IO) time is estimated at 10ms (i.e., 20/2039), while the computation time is a marginal 0.06ms (i.e., 20/312,000). These calculations underscore that the IO time, rather than the computation time which is contingent on the hardware’s FLOPs capacity, is the predominant factor influencing the overall inference latency of LLMs, due to its strong correlation with the model size and the available memory bandwidth.
@@
Various techniques, such as quantization (Lin et al., 2023; Frantar et al., 2022), sparsity (Lu et al., 2023; Zhou et al., 2021), pruning (Zhang et al., 2022b; Lagunas et al., 2021), distilling (Jeon et al., 2023; Hsieh et al., 2023), and tensor decomposition (Ma et al., 2019; Wang et al., 2023), have been proposed to reduce the LLMs’ size and the IO consumption time for predicting each token in LLMs. However, these techniques have been found to result in a degradation of accuracy. To address the challenge of predicting more tokens within a fixed IO consumption time, non-autoregressive neural machine translation (NAT) (Gu et al., 2017) and iterative parallel decoding (Santilli et al., 2023) have been introduced and successfully applied in translation tasks. Unfortunately, this approach has shown limited effectiveness in question-answering scenarios.
Recently, speculative decoding with a draft model has become a popular strategy. However, this strategy necessitates extra training efforts (Stern et al., 2018; Cai et al., 2024; Bae et al., 2023) or a smaller auxiliary model capable of producing drafts (Xia et al., 2023; Chen et al., 2023; Miao et al., 2024), and they may worse the memory burden with additional parameters or models. In light of this, training and assist-model free strategies are proposed, such as LLMA (Yang et al., 2023a) and LookaheadDecoding (Fu et al., 2023). The LLMA algorithm relies on a text-matching technique, which is effective within document retrieval domains, tends to underperform in other applications. LookaheadDecoding, on the other hand, incorporates the Jacobi iteration for draft generation, which, despite its innovative design, may face computational bottlenecks that impair its overall effectiveness.
@@
Table 1 summarizes the acceleration techniques discussed previously. To address their limitations, we introduce Lookahead, a groundbreaking framework that incorporates a trie-tree-based retrieving strategy and a multi-branch-based parallel Verification and Accept (VA) strategy. Lookahead leverages a trie tree to record the n-gram tokens of input prompts and generated responses. The draft is retrieved based on the provided context tokens, allowing for extremely fast draft generation. The draft tokens arranged in a logical tree structure are efficiently processed in the Verification and Accept (VA) process. Furthermore, we have implemented an adaptive strategy to optimize the retrieval process, effectively striking a balance between memory and computation requirements. It shows that Lookahead proves its superior performance in accelerating LLMs’ inference, compared with the existing state-of-the-art (SOTA) acceleration method.
@@
Methods Accuracy- Training- Multi- Low-cost-
lossless or-Assis- branch- draft-
Model-free draft generation
Quantization (Lin et al., 2023; Frantar et al., 2022) $\times$ - - -
Sparsity (Lu et al., 2023; Zhou et al., 2021) $\times$ - - -
Pruning (Zhang et al., 2022b; Lagunas et al., 2021) $\times$ - - -
Distilling (Jeon et al., 2023; Hsieh et al., 2023) $\times$ $\times$ - -
Tensor-decomp (Ma et al., 2019; Wang et al., 2023) $\times$ $\times$ - -
Early-exit (Li et al., 2021; Xin et al., 2020) $\times$ $\times$ - -
Block decoding (Stern et al., 2018) $\times$ $\times$
Spec decoding (Xia et al., 2023) $\times$ $\times$ $\times$
SpecInfer (Miao et al., 2024) $\times$ $\times$
FREE (Bae et al., 2023) $\times$ $\times$
LLMA (Yang et al., 2023a) $\times$
LookaheadDecoding (Fu et al., 2023) $\times$
Lookahead (ours)
@@
Moreover, to gain a wide range of applications, we have implemented our framework based on the transformers library of Hugging face [^1] https://huggingface.co/, by extending a generation mode named lookahead generation, which supports the greedy search and sample generation strategy. We have also currently applied Lookahead to the most recent LLMs, such as GLM(Du et al., 2021), Llama(Touvron et al., 2023), OPT(Zhang et al., 2022a), GPT2(Radford et al., 2019), BLOOM(Scao et al., 2022), ChatGLM(Zeng et al., 2022), Baichuan(Yang et al., 2023b) and Qwen(Bai et al., 2023), InternLM(Team, 2023), Mistral(Jiang et al., 2023), Mixtral MoE(Jiang et al., 2024), etc. The aforementioned models can be easily adapted to integrate Lookahead, our well-designed framework, with only minor code modifications of approximately 20 lines, which can be found in our repository.
@@
  • We empirically quantify that the main bottleneck of LLM inference is the IO bandwidth, rather than the computation bound.
@@
  • We elaborately adapt Lookahead to the most recent LLMs without any assistance of smaller models and have released our work with open source.
@@ 2. Related Work
Recently, several strategies have been proposed and developed to enhance the inference speed of LLMs while maintaining the output quality within an acceptable range. One such strategy is the non-auto-regressive approach, specifically non-auto-regressive translation (NAT) (Gu et al., 2017), primarily employed in translation tasks (Kasai et al., 2021; Huang et al., 2021; Saharia et al., 2020). However, it is essential to note that there are significant distinctions between translation tasks and general language model (LLM) scenarios, which may lead to subpar performance when applying the NAT strategy to LLM decoding.
@@
To address this limitation, Huang et al. (Huang et al., 2021) introduce a layer-wise iterative method wherein each layer utilizes the decoding results and embeddings from the preceding layers. This approach involves training each layer through maximum likelihood estimation to predict the outcomes of subsequent decoding layers. On the other hand, Santilli et al. (Santilli et al., 2023) formalized the standard greedy auto-regressive decoding strategy by employing a parallel Jacobi and Gauss-Seidel fixed-point iteration. It initializes the next tokens using special tokens and performs iterative decoding until convergence. However, all these methods may suffer from the risk of accuracy degeneration, since the manipulated model deviates from its original version.
@@
Recently, there has been a proposal for accuracy-lossless acceleration to enhance the auto-regressive decoding strategy as illustrated in Table 1. In this approach, a block-wise parallel decoding strategy was introduced by (Stern et al., 2018). In this strategy, each subsequent token is independently and parallel predicted as a proposal using an additional transformer model, which consists of a multi-output feed-forward layer and can be fine-tuned or distilled for optimal performance. Then the proposals are directly compared against the output tokens generated by the original decoder. The longest verified tokens are then selected for acceptance as the current step’s output. However, frequent failures during the verification process may be occur due to its reliance on a singular predictive branch. To overcome this drawback, Medusa (Cai et al., 2024) employs multiple heads to simultaneously predict drafts, enhancing the robustness of the process. On the other hand, FREE (Bae et al., 2023), uses the shallow layers of a model to generate drafts, instead of the final layer. to address this dependency. It proposals a synchronized parallel decoding strategy to ensure accuracy without loss. Despite this strategy’s remarkable acceleration, more efforts are needed to train the extra layer.
To address the aforementioned issue, speculative decoding has been proposed (Xia et al., 2023; Chen et al., 2023; Miao et al., 2024; Zhang et al., 2024; Hong et al., 2024). These works utilize a smaller model as a draft predictor. For instance, the Bloom 7.1B model can be employed as a draft model for the larger 176B model. However, the proposed works above may face significant practical challenges. Primarily, the availability of a smaller-scale model, typically one-tenth the size of the larger model, is not always feasible for a range of model series, such as Llama-7B (Touvron et al., 2023) and ChatGLM-6B (Zeng et al., 2022). Furthermore, larger models are often fine-tuned to specific applications, necessitating a parallel fine-tuning of the smaller helper model to produce comparable drafts. This requirement complicates the deployment and diminishes the system’s accessibility.
@@
To overcome this challenge, model-free prediction strategies have been introduced to achieve accurate predictions without relying on a specific model. One such strategy is presented in Ge et al. (Ge et al., 2022), which utilizes an input-guided method that copies content from the input sentence through prefix matching. Another strategy LLMA, proposed by Yang et al. (Yang et al., 2023a), employs a prefix matching approach to retrieve content from either the input sentence or a document database. However, it is worth noting that the model-free prediction strategies mentioned above utilize tokens in a single draft manner, failing to fully utilize the GPUs. Recently, another training-free and assist-model-free method named LookaheadDecoding (Fu et al., 2023) explores the multi-branch strategy with employing Jacobi iteration and speculative decoding simultaneously. Nonetheless, this method incurs a substantial overhead associated with the generation of drafts, which consequently attenuates the potential for acceleration.
@@ 3.3. Singe-Branch Strategy
Several most recent methodologies (Stern et al., 2018) (Yang et al., 2023a) have been proposed to generate a sequence of tokens at each decoding step, with the purpose of promising a higher l to accelerate the LLMs’ inference speed V. In these works, according to the prompt and the output tokens at the previous decoding step, a branch of tokens, named single-branch draft, have been obtained through the small model or the document reference, and efficiently validated by running the LLM in a single forward process.
Despite the single-branch strategy’s success in accelerating LLMs’ inference, the generated tokens’ length per decoding step, namely l, cannot be guaranteed as long as we wish. The previous works (Stern et al., 2018) (Yang et al., 2023a) empirically conclude that there is an upper limit for l and the inference speed V.
@@ 3.4. Discovering GPU’s FLOPs Redundancy
To explore how far EDL can be extended at each decoding step, we conduct a case study through discussing the decoding length’s impact on the overall consuming time of LLMs’ single forward process, whose result can be found in Figure 1. It should be noted that we apply AntGLM-10B (which is trained from scratch with the GLM structure (Du et al., 2021; Zeng et al., 2022)) model with single Nvidia A100 GPU as an example.
Figure 1. Decoding length’s impact on the overall consuming time of LLMs’ single forward process. Even the forward FLOPs is linear to the decoding length,
In Figure 1, it is evident that as the decoding length increases in the LLM’s single forward process, the time consumption remains relatively constant in the initial stage, given a fixed context length. This observation can be attributed to several factors. Firstly, certain time requirements, such as kernel launching, operator loading, reading, and verification, are fixed and dependent on the GPU bandwidth. These factors are unrelated to the computational complexity or decoding length. Secondly, when the decoding length is relatively short, the resulting small matrix blocks may not take full advantage of CUDA and Tensor cores, indicating that GPU shows its FLOPs redundancy in this scenario. Consequently, the differences in overall time consumption between decoding lengths of 16 and 128 in the LLM’s single forward process is negligible, provided the context length is 256. This finding supports the concept that the decoding time per decoding step, represented as t(l), remains nearly constant within a specific range of generated token lengths l. However, when the decoding length exceeds 128 while the context length is 256 in the LLM’s single forward process, the larger matrix block size introduces a more complex calculation logic for the tensor cores. Consequently, the overall inference time shifts to the second stage, leading to a gradual increase in overall inference time.
@@
It can be clearly seen that there is still a ”sufficient gap” between EDLs empirically concluded previously (Stern et al., 2018) (Yang et al., 2023a) and the CDL illustrated in Figure 1. In this scenario, we come up with two questions:
@@ 4.1. Overview of Lookahead
In contrast to the single-branch strategy such as (Yang et al., 2023a), which only considers one available draft without considering the existence of other drafts, the multi-branch strategy retrieves multiple drafts simultaneously. These drafts are then efficiently decoded and validated in parallel through the Verification and Accept (VA) process, progressively. The VA process then identifies the correct sub-sequence for each draft and retains the longest sub-sequence as the output tokens. Figure 2 presents the progress overview including various draft retrieving and verifying strategies.